OpenStudy (anonymous):

The local beach resort is offering two weekand specials. One special includes a 2-night stay with 3 meals which costs $195. The other special includes a 3-night stay with 5 meals and costs $300. What is the cost of a 1-night stay? a. $15.00 b. $35.00 c. $55.00 d. $75.00

OpenStudy (amistre64):

we have 2 variables and 2 equations to set up

OpenStudy (amistre64):

2n + 3m = 195 3n + 5m = 300 ; now we can solve for n or m or both

OpenStudy (anonymous):

So what's the answer?

OpenStudy (amistre64):

the answer is to work thru the problem :)

OpenStudy (amistre64):

we can solve it by elimination prolly the easiest

OpenStudy (amistre64):

nah, this way is fine

OpenStudy (anonymous):

multiply the 1st equation with 3 and 2nd equation with 2

OpenStudy (anonymous):

\[6n=390-9m\]\[6n=600-10m\]

OpenStudy (anonymous):

I just need the answer to see if my work was right

OpenStudy (anonymous):

the answer is d.

OpenStudy (amistre64):

im sure you do, but to keep the place from not being used for unethical purposes, we really should help you thru the problem instead of simply giving out answers ...

OpenStudy (anonymous):

I knew my answer was right.

OpenStudy (amistre64):

:) good, what was your answer?

OpenStudy (anonymous):

My answer was D.

OpenStudy (anonymous):

u r right

OpenStudy (anonymous):

good , glad i could help

OpenStudy (amistre64):

2(75) + 3m = 195 150 + 3m = 195 3m = 195-150 m = 45/3 m = 15 3(75) + 5m = 300 225 + 5m = 300 5m = 300-225 m = 75/5 m = 15 i agree :)